Bombarding a bismuth-209 (Bi-209) target with a beam of another nuclide produces bohrium-262 (Bh-262) and a neutron. Identify the isotope used in the bombardment. (in order in order to identify the nuclide used in the bombardment, start with writing the balanced equation for this nuclear reaction)
Answer #1

Let the reactant be X

The decay eqn is:
₈₃²⁰⁹Bi + X -> ₁₀₇²⁶²Bh + ₀¹n



Balance mass number on both sides
mass of Bi + mass of X = mass of Bh + mass of n
1*209 + M = 1*262 + 1*1
M = 54

Balance atomic number on both sides
atomic of Bi + atomic of X = atomic of Bh + atomic of n
1*83 + A = 1*107 + 1*0
A = 24
Atomic number 24 is for element Cr
So, the reactant is
54-Cr
Answer: 5424Cr